OpenStudy (anonymous):

Find the sum, if it exists. 75+15+3+

OpenStudy (diyadiya):

you want to add it ? o.O

OpenStudy (diyadiya):

75+15+3+ ?

OpenStudy (anonymous):

sorry the problem is written as 75+15+3+... Not sure what to do?

OpenStudy (zepp):

This a geometric serie. 75,15,3.... Common ratio being 3, first term 75 And add up to infinite.. :|

OpenStudy (anonymous):

That's a geometric series with r = 1/5. Use the summation formula :-)

OpenStudy (zepp):

Common ratio 0.2 sorry

OpenStudy (anonymous):

It converges, r < 1.

OpenStudy (anonymous):

In general: \[ S_{\infty} = \frac{a_1}{1-r} \]for |r| < 1.

OpenStudy (anonymous):

Plug in now @onley , what did you get?